OpenStudy (anonymous):

What is 3^2/3 equal to? A.3√9 B.2√9 C.3√27 D.2√27

OpenStudy (margotlakus):

I'm at school myself and bored in class so I can't help as much as I'd like... :c BUT try typing the answers into your calculator and hitting the button. hope that helps! c:

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

@alakara_kibo :)

OpenStudy (anonymous):

give me a sec :)

OpenStudy (anonymous):

its A :)

OpenStudy (anonymous):

3√5 Thanks :) take a look at this one: Simplify ----- 4√5 A.5^1/4 B.5^1/12 C.5^7/12 D.5^4/3

OpenStudy (anonymous):

whoops it was 2/27 for the last one had to fix an error and ok :)

OpenStudy (anonymous):

Oh ok! hehe thank you :P

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

u can always ask me question :) i will be glad to help:)

OpenStudy (anonymous):

Aw thanks! I really appreciate your help. I have like ten thousand more questions lol, but I'll just ask a few :P

OpenStudy (anonymous):

its alright if u ask alot i will be glad to help cuz i like math and is ahead of my own class :) so ur fine :)

OpenStudy (anonymous):

Awesome, thanks! Well this one looks really complicated :S The expression 1000(1.002306)^4x shows the quarterly interest rate Patty earns after x years. Which expression shows the x isolated in the exponent in an equivalent form? A.1000(0.250576)^x B.1000(1.009256)^x C.1000(4.009224)^x D.1000(1.000576)^x

OpenStudy (anonymous):

its C C.1000(4.009224)^x

OpenStudy (anonymous):

Ok! mmm what about this one: A function is shown below: f(x)=(0.07)^x What does the function represent? A.exponential growth of 7% B.exponential decay of 7% C.exponential growth of 93% D.exponential decay of 93%

OpenStudy (anonymous):

B is the best answer:)

OpenStudy (anonymous):

Ok, thanks, you're like an angel!!!! :D... I'll ask one more and then I have to go to a class but I'll be back in an hour hehe. The graph of f(x)=(0.5)^x is replaced by the graph of g(x) = (0.5)^x - k. If g(x) is obtained by shifting f(x) down by 3 units, the value of k is ______

OpenStudy (anonymous):

k= 1.5

OpenStudy (anonymous):

Hey! I'm back :)... is it ok if I ask you a few more?

OpenStudy (anonymous):

sure im also helping one other and that person has 12 questions for me to answer so i can help

OpenStudy (anonymous):

Oh!! ok thanks! well look at this one: the following function represents the value of a house after x years f(x) = 242,000(0.96)^x What does 0.96 represent? A.the present value of the house B.the value of the house after x years C.the increase in the value of the house per year, which is 4% D.the decrease in the value of the house per year, which is 4%

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Ok, what is the value of y in the solution to the following system of equations? 2x+y=-4 5x+3y=-6

OpenStudy (anonymous):

y= 1.42

OpenStudy (anonymous):

ok, Determine the type of boundary line and shading for the graph of the inequality 3x+y> 10 (the dash beneath the > means greater than or equal to :) ) - A.dashed line with shading on the side that includes the origin B.solid line with shading on the side that does not include the origin C.dashed line with shading on the side that does not include the origin D.solid line with shading on the side that includes the origin (I think is B, but I'm not sure)

OpenStudy (anonymous):

there is a dash beneath the > I just don't know how to put it lol ._.

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

and i know wat u meant lol

OpenStudy (anonymous):

hehe ok thank you. choose the system of equations that matches the following graph A. 3x+4y=24 3x+y=-3 B. 3x+4y=24 3x-y=3 C. 3x-4y=-24 3x+y=-3 D. 3x-4y=-24 3x-y=23

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Ok :P... Well there two last questions I'm really stuck at, but here you go: Describe the graph that best represents the function f(x)=4(1.5) (there are no options ._.)

OpenStudy (anonymous):

Oh wait! there are options! it's just that my browser wasn't loading them lol, let me show them to you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just let me know when u post it :)

OpenStudy (anonymous):

Here you go :)

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

Awesome, I think I'm almost done with the ones I don't know! I have like four more I really have no idea lol

OpenStudy (anonymous):

alright well it will talk me a few to reply back cuz im sharing my comp but i will help as much as i can ok

OpenStudy (anonymous):

Alright! thank you very much! I attached the table I'm having troubles with!

OpenStudy (anonymous):

D is ur answer:)

OpenStudy (anonymous):

Thanks! what about this other table?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

This one?

OpenStudy (anonymous):

aaand the last one!

OpenStudy (anonymous):

ok sorry its talking me so long to write back cuz my comp crashed on me:(

OpenStudy (anonymous):

the first one is A and the second one is D